# Decide less, communicate more: On the construct validity of end-to-end fact-checking in medicine

### v2

@@ 6 Conclusion
Appendix A All Claims from Last Split
@@
Post (r/ADHD): Hello, menstruating people! How do your cycle and ADHD influence each other and how do you deal with it?
EDIT: After getting your responses I am reflecting again how medicine does not give a shit about women. It’s truly insane. Thank you! Hello! I have never paid too much attention to my menstrual cycle since it was never particularly bothersome. Now that I take methylo I feel big changes in how I function during the cycle. Like last 10 days of the cycle, my medication kind of stops working… That is like 1/3 of the time. I know it’s still better than without meds nevertheless, it makes establishing a routine quite challenging. My doc suggested trying contraceptive pills, but I am not even sexually active ATM so taking more medication, with potential side effects, does not excite me. I know there are herbs that are proven to be helping with regulating the cycle but I don’t know if they would help with ADHD symptoms? Any tips?
Population: People with ADHD,
Intervention: Herbs,
@@
Post (r/ADHD): Stimulants vs. Sodium
Im wondering if anyone else has experienced this. I find that my stimulant medications (Adderall IR and Vyvanse) make me very sensitive to salt. If I have a higher sodium meal (eg ramen or canned soup, or even just mustard on my sandwich), I get very bloated. Its uncomfortable and lasts for a few days. Whenever I take a break from my meds, this doesnt happen. Ive had labs done for it in the past and it doesnt seem like anything medically problematic, but its uncomfortable and it really stresses me out.
Population: People with ADHD,
Intervention: Stimulant medications (Adderall IR and Vyvanse),
@@
Population: Patients with Cystic Fibrosis (implied by the subreddit r/CysticFibrosis),
Intervention: Trikafta (a medication),
@@
Post (r/CysticFibrosis): Anyone with sinus issues drinking pineapple juice?
It’s a weird question, but I saw a post about pineapple juice being good for sinus issues (helps with the inflammation) and just wondered if anyone has done this? Some people were commenting about the high sugar content in pineapple juice not being good, but they get around that by taking a supplement instead of drinking the juice. Anyone?
Population: Patients With Cystic Fibrosis,
Intervention: Pineapple Juice,
@@
Post (r/CysticFibrosis): Trikafta & PMDD
So, I believe trikafta has given me PMDD premenstrual dysphagia disorder. Every month, the week before my period I have extreme anxiety in a running dialouge in my head that is constantly negative. I’ve never been this way before. I also have horrible hormonal acne on my back & forehead which are very new to me as well. My question is: any one else having this problem? My Dr said they are noticing a "negative interaction with estrogen and trikafta". Anyone find anything that helps??
Population: Patients with cystic fibrosis (implied by the Reddit thread r/CysticFibrosis), specifically females of reproductive age,
Intervention: Trikafta,
@@
Table 7: All claims from the last split. Claim in the post is highlighted. Given population annotation is highlighted in blue. Intervention is highlighted in pink. Outcome is highlighted in green.
We present all the claims from the last split of annotations we collected from medical experts in Table 7.
Appendix B Pilot Inter-evaluator Agreement
The initial agreement for the pilot set of 10 claims is shown Table 8.
Type $\kappa$
Population 0.379
Intervention 0.355
Outcome 0.279
Overall 0.347
Tab Support 0.353
Overall Support 0.191
Table 8: Inter-evaluator agreement measured through Fleiss’ $\kappa$ for the training set of 10 claims.
Appendix C Refined Guidelines Agreement
The agreement for the refined guidelines is shown in Table 9.
Type $\kappa$
Population 0.420
Intervention 0.131
Outcome 0.173
Overall 0.161
Tab Support 0.207
Overall Support -0.011
Expert Support -0.120
Table 9: Inter-evaluator agreement measured through Fleiss’ $\kappa$ for the updated guidelines test set of 5 claims.
Appendix D Why we Added PIO Elements?
After our pilot and refinement set, we realized that a source of disagreement in the annotators was because of the lack of PIO elements to focus on.
There was sometimes disagreement in cases where there are multiple PIO elements, leaving it to the discretion of the medical experts creates disagreement, as some choose to consider both the PIO elements or just focus on one. An example is shown in LABEL:table:prednisone_ex.
There are rare instances also where the annotated PIO highlights are incorrect, for example, accidentally labeling a population as an intervention. Therefore, we devise a LLM-based pipeline to extract the Population, Intervention, and Outcome elements from the post. Our prompt and expert validation are detailed in Appendix E.
Table 10: This figure demonstrates how experts can misinterpret the claim to evaluate without explicit guidance on which PIO elements to focus on in the claim.
@@ Appendix D Why we Added PIO Elements?
Post (r/lupus): Cytoxan and prednisone
Rheumatologist says cellcept failed to protect my kidneys and now I have developed lupus nephritis.Im so upset. Prednisone messed up my hips so badly that they both need to be replacedI dont want to get back on it but rheumatologist says its to bring the inflammation down in my kidneys. Ive never been on Cytoxan but the side effects sound identical to a lupus flare. How am I supposed to be positive with news like this? I feel so defeatedI dont know what to do.
@@
Appendix E PIO Extraction
We used automatic method of LLAMA 3.1 405B.
After preliminary prompt testing and consultation with our medical expert, we utilized PIO element definitions and questions in our prompt from Duke https://guides.mclibrary.duke.edu/ebm/pico:
PATIENT OR PROBLEM
INTERVENTION, EXPOSURE, PROGNOSTIC FACTOR
OUTCOME
Text posted by someone in Reddit thread r/[sub][sub_description]:
[post]
Highlighted claim:
[claim]
On a sample set of 55 samples (random 5 claims sampled per each of the 11 conditions), our medical expert from an esteemd hospital reviewed and reported % accuracy on the PIO extractions. He said it was ready to go, except for a few minor comments including the implied message as well as there were a few instances where the extracted Population and Intervention overlapped. He reported accuracy of above 90% for each element.
RCT Verifiability of the PIO elements doesn’t necessarily mean RCT verifiability of the claim.
This was our validation for our PIO extraction and we used it to select claims and give PIO elements.
Appendix F Retrieval Strategies
For the first phase of this research of claims and abstracts we evaluated for the pilot, we used state-of-the-art embedding model, Alibaba-NLP/gte-large-en-v1.5, which was the top performing open-sourced embedding model with lightweight parameters on 06/03/2024 according to the HuggingFace MTEB Benchmark, which ensured feasible use to embed a database of 800,000 RCT abstracts with our limited resources.
For the next phase, we did more extensive experiments with dunzhang/stella_en_400M_v5 evaluating different prompts and setups for retrieval.
We used SOTA embedding model, dunzhang/stella_en_400M_v5, which was the SOTA open-sourced embedding model with lightweight parameters on 09/11/2024 according to the HuggingFace MTEB Benchmark, feasible to use to embed a database of 800,000 RCT abstracts with our limited resources.
The overall scores are explained in Table 11
The claims are shown in LABEL:table:retrievalclaims
Strategy Query Document Pop. Inter. Out. Overall
S2P Question PIO PIO 10.9 8.1 9.8 7.4
S2P PIO Abstract 11.3 8.7 10.6 8.3
S2P PIO PIO 11.4 8.6 10.3 8.1
S2S PIO Abstract 11.8 8.4 11.2 8.3
S2S PIO PIO & Abstract 11.1 8.1 10.2 7.9
S2S PIO PIO 11.5 7.8 9.6 7.6
S2S PIO PIO & Title & Abstract 11.2 8.2 10.2 7.8
S2S PIO Title & Abstract 11.6 8.3 10.6 8.0
Table 11: Retrieval Strategies Results on Test Set of 5 Claims. 15 is full score: each claim is averaged across 10 abstracts. The score of 1 is irrelevant, 2 is somewhat relevant, and 3 is relevant.
Post P I C All
Grapefruit and siezures
Post (r/Epilepsy): Grapefruit and siezures
Guys, I am wondering if you have any issues or know about interactions between oxcarbazepine and/or levetiracetam and grapefruit? I believe it may make those medications work differently, but I am not sure.
Population: Epilepsy patients (implied by the subreddit r/Epilepsy and the mention of seizure medications),
Intervention: Grapefruit consumption (in interaction with oxcarbazepine and/or levetiracetam),
Outcome: Medication efficacy (i.e., how the medications work)
20.7 11.1 19.3 10.4
COVID-19
Post (r/MultipleSclerosis): Anyone take Paxlovid?
Hi everyone - one of my kids just tested positive for Covid so assuming I will as well in the next few days. I contacted my PCP and neuro who said that as soon as I test positive they would recommend getting me on Paxlovid as long as we can find it (vs monoclonal antibody infusions, as apparently those are not working quite as effectively against Omicron). Ive been doing my research on the drug but am curious if anyone has taken it yet. What was your experience? The worst side effect she mentioned was like a battery acid taste in your mouth for the entire 5-day regimen. Sounds gross but doable. Thanks in advance for any insight.
Population: COVID patients (specifically those with Omicron variant, and implied to be similar to the author who has Multiple Sclerosis) ,
Intervention: Monoclonal antibody infusions (vs Paxlovid, but the claim specifically refers to the effectiveness of monoclonal antibody infusions) ,
Outcome: Effectiveness (implied to be against Omicron variant, but not a specific measurable outcome)
23.1 18.1 22.5 18.1
Balloon Sinuplasty
Post (r/Sinusitis): After your balloon sinuplasty, how long did it take you to get rid of your face pressure and shallow breathing? Its been 5 days since my procedure and I still have facial pressure. This effects me in negative ways like not being able to focus with my eyes. I also have trouble automatically deep breathing. Im just curious how long does it typically take to recover from these symptoms? I was told I should experiencing some relief after 2 days but still nothing. Im starting to wonder if this is dental related.
Population: Patients who have undergone balloon sinuplasty (similar to the author of the post, who has sinusitis),
Intervention: Balloon sinuplasty,
Outcome: Relief from facial pressure and breathing symptoms (specifically within 2 days after the procedure)
8.3 8.2 8.7 7.9
Elvanse
Post (r/ADHD): Picking the correct meds? So I’m UK based and have just started titration through PsychiatryUK after waiting nearly a year.
Started on Xaggitin XL 18mg ( now on 36mg and then will go to 54mg ). It seems to work … I wake up a lot easier as before I would wake up groggy and just completely unfunctional .. which would then lead to me being easily distracted etc, and my focus throughout the day seems improved … mood seems improved … but, at around 4pm I crash hard and just feel exhausted at night. I had some family over at the weekend and was unable to keep up in conversation. Also after a few days it seems to have less of an effect … but then I up the dose and it works again. Obviously this won’t be sustainable and I will eventually reach max dose. I would like to try Elvanse next as I’ve heard people say the come down is less brutal … but it takes longer to kick in, which is a problem for me as I find it most difficult in the morning. I could wake earlier and take it, and then go back to bed … or I’ve also seen some people who take Amfexa as a top up in the morning and then take Elvanse for the rest of the day. This seems totally ideal for me, and has the benefit where I can forgo the meds on weekends if I’m not working / doing much - but have the quick release to fall back on in the event something crops up. The question is - can I request this? I scared to accept a med just because it works somewhat, as changing it seems like it would be a pain in the ass .. so I’d like to make an informed decision.. but also don’t want to waste their time. I’m also scared to even mention Amfexa for fear of being labeled as someone just wanting a high.
Population: ADHD patients (specifically, those similar to the author who experience a crash with their current medication) ,
Intervention: Elvanse,
Outcome: Severity of the "come down" (i.e., the crash or comedown experienced by the patient after the medication wears off)
18.4 15.7 14.6 13.1
Probiotics IBS
Post (r/ibs): Some Info on Probiotics for IBS
So I posted this in a comment and it was suggested that I make this info its own post. I did a bunch of research into what probiotics have shown to be effective for IBS in various research studies. From the links below, the collective summary is that these probiotic strains (alone or in combination with others) were effective in one or more studies in reducing one or more IBS symptoms in frequency and/or severity: Saccharomyces Boulardii CNCM I-745 (actually a yeast), Bifidobacterium longum 35624, Bifidobacterium infantis, Lactobacillus acidophilus, Lactobacillus plantarum 299v, Lactobacillus rhamnosus GG, Saccharomyces cerevisiae CNCM I-3856, Bifidobacterium lactis HN019, Bifidobacterium lactis BB-12, Lactobacillus acidophilus NCFM, Bifidobacterium lactis Bi-07, and Bifidobacterium bifidum MIMBb75. I may have missed a few in the above list, but the main gist is that Bifidobacterium and Lactobacillus species probiotics, as well as Saccharomyces Boulardii, are the main researched probiotics that show some kind of improvement for IBS (though it also depends on the type of IBS, and most of them helped with one or some symptoms but not all of them - see links below for more detail). Also, many of the studies couldn’t draw conclusive evidence because there just wasn’t enough data/research on specific probiotics yet, or because the studies didn’t fulfill certain criteria. But they were able to state what evidence they did have, and identify what is likely to be the case based on findings so far. Bifidobacterium longum 35624 used to be called Bifidobacterium infantis 35624 and is still sometimes referred to that way. It is patented by Align, which is the brand name for that probiotic. Floraster is a leading brand name for Saccharomyces Boulardii CNCM I-745. There are other good probiotic brands too though (that have similar or different probiotic strains). A lot of this will be trial and error as you find what works for you. Be sure to give each probiotic or probiotic blend you try at least a 4 week trial, since it takes a bit of time for your body to adjust and for the probiotics to get to a certain level in your system to start working. It’s a good idea to make sure that the probiotics are third-party tested and made in a GMC (or equivalent) registerred facility, to be sure you are getting exactly what’s on the product label. I’m not sure exactly how it works in other countries, and that advice is US-based, but some level of quality assurance is generally advised. Also, for those who aren’t used to reading research papers like this, usually the important information is in the Results and/or Discussion section. Sometimes they put a summary of those near the top, and sometimes you have to scroll to the end of the article. It’s also a good idea to read the abstract/summary at the top of the article so you have an idea of what the paper is about. Reading the other sections of the paper will give you more detail into what they did for the study, which will give more context, but isn’t usually necessary to read the whole thing to understand the end results, which tell you what they found about the probiotics in relation to IBS. And of course, always talk to your doctor about medical stuff! I’m just providing this information in case it might be helpful. I am NOT giving medical advice or telling you all to take probiotics or endorsing any brands or types or anything. Whether you decide to take probiotics or not is up to you and your doctor. This is just an info dump in case anyone wanted to know this stuff. Here are the links: https://www.thelancet.com/journals/eclinm/article/PIIS2589-5370(21)00434-X/fulltext https://www.sciencedirect.com/science/article/pii/S1590865819309594 https://www.ncbi.nlm.nih.gov/labs/pmc/articles/PMC7279071/ https://www.optibacprobiotics.com/learning-lab/in-depth/gut-health/which-probiotics-are-best-for-ibs#IBS-U https://www.frontiersin.org/articles/10.3389/fphar.2020.00332/full#B68 https://badgut.org/information-centre/a-z-digestive-topics/probiotics-for-irritable-bowel-syndrome/ https://www.ncbi.nlm.nih.gov/labs/pmc/articles/PMC3296087/ https://pubmed.ncbi.nlm.nih.gov/33745570/ https://onlinelibrary.wiley.com/doi/abs/10.1111/eci.13201
Population: Patients with IBS (Irritable Bowel Syndrome),
Intervention: Probiotics (various strains),
Outcome: Reduction in IBS symptoms (frequency and/or severity)
20.3 13.1 17.4 13.9
Table 12: Retrieval Test Ratings (Population, Intervention, Outcome, Overall) by Claim. Full Score is 24. Claim in the post is highlighted. Given population annotation is highlighted in blue. Intervention is highlighted in pink. Outcome is highlighted in green.
Appendix G Annotation Guidelines
@@ Appendix H Annotator Recruitment
We recruited five of our experts from Upwork, a process that took four weeks. During this phase, we received 117 proposals on Upwork, and reached out to 19 people with a sample annotation task for claim in 13 to gauge their instruction-following capability and and their medical expertise through the quality of their annotation and plain language explanations. From this, we selected 7 of the most qualified candidates by explanation quality, interviewed them for final fit, and ultimately chose 5 which best met our expectations. Our experts worked anywhere between 3 to 20 hours per week of annotation. Our medical experts on average took 20 minutes to annotate each claim end-to-end. We paid our experts a range between \$22 and \$35 through Upwork.
@@
Table 13: The claim is highlighted. The RedHOT intervention annotation is highlighted in pink.
Our total spend for the annotation study was \$1432.20.
Appendix I Pilot Claims

### v3

@@ Acknowledgments
Appendix A Trikafta Claim Example
Dandruff and Trikafta
Population: Patients with Cystic Fibrosis (implied by the subreddit r/CysticFibrosis)
Intervention: Trikafta (a medication)
Outcome: Flaky scalp or dandruff
Appendix B All Claims from Final Split
@@
Post (r/ADHD): Hello, menstruating people! How do your cycle and ADHD influence each other and how do you deal with it? EDIT: After getting your responses I am reflecting again how medicine does not give a shit about women. It’s truly insane. Thank you! Hello! I have never paid too much attention to my menstrual cycle since it was never particularly bothersome. Now that I take methylo I feel big changes in how I function during the cycle. Like last 10 days of the cycle, my medication kind of stops working… That is like 1/3 of the time. I know it’s still better than without meds nevertheless, it makes establishing a routine quite challenging. My doc suggested trying contraceptive pills, but I am not even sexually active ATM so taking more medication, with potential side effects, does not excite me. I know there are herbs that are proven to be helping with regulating the cycle but I don’t know if they would help with ADHD symptoms? Any tips?
Population: People with ADHD
Intervention: Herbs
@@
Post (r/ADHD): Stimulants vs. Sodium Im wondering if anyone else has experienced this. I find that my stimulant medications (Adderall IR and Vyvanse) make me very sensitive to salt. If I have a higher sodium meal (eg ramen or canned soup, or even just mustard on my sandwich), I get very bloated. Its uncomfortable and lasts for a few days. Whenever I take a break from my meds, this doesnt happen. Ive had labs done for it in the past and it doesnt seem like anything medically problematic, but its uncomfortable and it really stresses me out.
Population: People with ADHD
Intervention: Stimulant medications (Adderall IR and Vyvanse)
@@
Population: Patients with Cystic Fibrosis (implied by the subreddit r/CysticFibrosis)
Intervention: Trikafta (a medication)
@@
Post (r/CysticFibrosis): Anyone with sinus issues drinking pineapple juice? It’s a weird question, but I saw a post about pineapple juice being good for sinus issues (helps with the inflammation) and just wondered if anyone has done this? Some people were commenting about the high sugar content in pineapple juice not being good, but they get around that by taking a supplement instead of drinking the juice. Anyone?
Population: Patients With Cystic Fibrosis
Intervention: Pineapple Juice
@@
Post (r/CysticFibrosis): Trikafta & PMDD So, I believe trikafta has given me PMDD premenstrual dysphagia disorder. Every month, the week before my period I have extreme anxiety in a running dialouge in my head that is constantly negative. I’ve never been this way before. I also have horrible hormonal acne on my back & forehead which are very new to me as well. My question is: any one else having this problem? My Dr said they are noticing a "negative interaction with estrogen and trikafta". Anyone find anything that helps??
Population: Patients with cystic fibrosis (implied by the Reddit thread r/CysticFibrosis), specifically females of reproductive age
Intervention: Trikafta
@@
Table 6: All claims from the final annotation split. The extracted claim span is highlighted in yellow; Population annotations are highlighted in blue, Intervention in pink, and Outcome in green.
Appendix C Rationale for Re-extracting PIO Elements
During the pilot and refinement rounds, we identified a key source of expert disagreement arising from underspecified focus when multiple Population, Intervention, and Outcome (PIO) elements were present in a claim. In such cases, experts differed on whether to consider all PIO elements or prioritize a subset, leading to inconsistent judgments. An illustrative example is shown in Table 7.
Although RedHOT provides PIO highlights for each claim, we observed occasional errors and inconsistencies in these annotations (e.g., mislabeling a Population as an Intervention), which could further contribute to ambiguity during expert review. To reduce this source of disagreement and standardize expert attention, we re-extracted explicit PIO elements from each post using an LLM-based pipeline, with prompt design and expert validation detailed in Appendix I.
@@ Appendix C Rationale for Re-extracting PIO Elements
Post (r/lupus): Cytoxan and prednisone Rheumatologist says cellcept failed to protect my kidneys and now I have developed lupus nephritis.Im so upset. Prednisone messed up my hips so badly that they both need to be replacedI dont want to get back on it but rheumatologist says its to bring the inflammation down in my kidneys. Ive never been on Cytoxan but the side effects sound identical to a lupus flare. How am I supposed to be positive with news like this? I feel so defeatedI dont know what to do.
@@
Table 7: Example illustrating expert disagreement when multiple PIO elements are present and no explicit guidance is provided on which elements to prioritize.
Appendix D Changes between annotation rounds
In the first round, ten claims were annotated without explicit PIO contextualization or the expert support field. In the second round, we refined the annotation guidelines to clarify label definitions; however, agreement on the refinement set of five claims remained low. Based on expert feedback, we substantially revised the setup for the third round by filtering out non–RCT-verifiable claims, re-extracting and providing PIO elements to guide expert focus, improving the retrieval system, and further clarifying the annotation guidelines. Experts then annotated five claims under this updated setting.
Appendix E Annotation Guidelines
@@
Appendix F Plain Language Explanation Guideline
  • (Optional) Medical Addendum at end.
Appendix G Annotation Interface
We used a web-based annotation interface to collect annotations from our expert medical annotators. Figure 3 shows how we present a claim with its surrounding context, extracted PIO elements for that claim, and retrieved abstracts corresponding to that claim. In this setup, we highlight the claim within the post in which its found along with any PIO spans as determined by data found in the RedHOT dataset. We also present extracted PIO elements (see Appendix I) in a separate box with a rewritten claim created by inputting these elements in a template. All of this information was provided for the benefit of the annotator to clearly understand the claim in question.
@@ Appendix H Annotator Recruitment
We recruited five medical experts via Upwork over a four-week period. We received 117 proposals, screened 19 candidates using a sample annotation task (Table 8) to assess instruction-following, medical knowledge, and explanation quality, and shortlisted seven candidates for interviews, from which five were selected. Experts worked between 3 and 20 hours per week and required approximately 20 minutes to annotate each claim end-to-end. Compensation ranged from \$22 to \$35 per hour. The total cost of the annotation study was \$1,432.20.
@@
Table 8:
Appendix I PIO Extraction
We used an automatic PIO extraction approach based on Llama-3.1-405B-Instruct. After preliminary prompt testing and consultation with a medical expert, we incorporated PIO element definitions and guiding questions adapted from Duke University’s PICO evidence-based medicine guidelines [^2] https://guides.mclibrary.duke.edu/ebm/pico.
PATIENT OR PROBLEM
INTERVENTION, EXPOSURE, PROGNOSTIC FACTOR
OUTCOME
Text posted by someone in Reddit thread r/[sub][sub_description]:
[post]
Highlighted claim:
[claim]
On a validation set of 55 samples (five claims randomly sampled from each of 11 conditions), a medical expert reviewed the extracted PIO elements. The expert reported over 90% accuracy for each element, noting only minor issues such as overlapping Population and Intervention spans and occasionally implied elements. Based on this review, the extraction pipeline was deemed sufficiently reliable for use in our study.
Appendix J Prompt for RCT Verifiability
Claim: [claim]
Text from r/[subreddit]: [post]
Appendix K Retrieval Configurations
We evaluated several retrieval configurations using the dunzhang/stella_en_400M_v5 embedding model on a small development set, with relevance judgments provided by a medical expert. Aggregate results are shown in Table 9. At the time of experimentation (09/11/2024), stella_en_400M_v5 was among the top-performing open-source embedding models on the HuggingFace MTEB benchmark while remaining computationally feasible for indexing approximately 800,000 RCT abstracts.
Strategy Query Document Pop. Inter. Out. Overall
S2P Question PIO PIO 2.2 1.6 2 1.5
S2P PIO Abstract 2.3 1.7 2.1 1.7
S2P PIO PIO 2.3 1.7 2.1 1.6
S2S PIO Abstract 2.4 1.7 2.2 1.7
S2S PIO PIO & Abstract 2.2 1.6 2.0 1.6
S2S PIO PIO 2.3 1.6 1.9 1.5
S2S PIO PIO & Title & Abstract 2.2 1.6 2.0 1.6
S2S PIO Title & Abstract 2.3 1.7 2.1 1.6
Table 9: Retrieval strategy results on a test set of five claims. Scores report average relevance per claim, computed over 10 retrieved abstracts (1 = irrelevant, 2 = somewhat relevant, 3 = relevant). The highlighted row indicates the configuration used in the main study.
Appendix L Implementation Considerations for the Communication Model
Appendix M Pilot Claims